# Co-Training Vision Language Models for Remote Sensing Multi-task Learning

### v1

@@ III-D Zoom-in Chain for UHR RS Images
Previous works on understanding UHR RS images, such as LRS-VQA [lrsvqa] and GeoLLaVA-8k [geollava8k], primarily focus on addressing the issue of excessive image tokens through visual token pruning. Although this approach has proven effective and computationally efficient, it typically requires additional training and is not well-suited for joint training with tasks using standard image resolutions.
@@
$~~~~~~~~~~\texttt{Assistant}:~{\color[rgb]{0,0,1}[x1,y1,x2,y2]}$
@@
$~~~~~~~~~~\texttt{Assistant}:~{\color[rgb]{0,0,1}\text{\textless Ground Truth\textgreater}}$ (5)
@@
To enable the model to learn zoom-in capabilities during training, we construct a specialized instruction tuning dataset for UHR RS image perception, named LRS-VQA-Zoom. The data pipeline is initiated by collecting three public, large-scale UHR RS datasets: DOTA1.0 [dota], GLH-Bridge [li2024learning], and STAR [li2024scene].
The methodology for generating the LRS-VQA-Zoom is extended from the pipeline in LRS-VQA [lrsvqa]. The final training corpus, totaling 302k samples, compromise three distinct subsets: 60k open-ended samples generated via rule-based templates, 159k open-ended samples synthesized using GPT-4V, and 83k samples in multi-choice-query form. Fig. 4 exhibits the examples from each subsets.
@@ III-E1 Response Normalization
During our extension of LMMRotate [lmmrotate], we observed a subtle yet important issue. In the LMMRotate, images without any objects were removed from the training set to improve efficiency, following common practices in conventional aerial detector. However, this approach can be detrimental when training a VLM, as encountering object-free images during inference often leads the model to hallucinate, producing false positive detections. To address this, RSCoVLM retains images without objects in the training process and explicitly trains the model to output “There is none.” for such cases, thereby mitigating hallucinations and improving detection reliability.
@@ III-E2 Evaluation Metrics
In conventional aerial detection tasks, mean average precision (mAP) is widely employed as the evaluation metric, requiring bounding boxes, class labels, and confidence scores for all detected objects. However, as discussed earlier, our model responses only include object categories and their corresponding spatial coordinates. We observed that the reliability of confidence estimation considerably influences the final mAP, which poses an inherent limitation for VLMs.
Figure 5 illustrates the influence of confidence scores on performance. The curves depict mAP values obtained when the confidence terms are replaced by either fixed or random constants, denoted as mAP with no conference scores, i.e., $\text{mAP}_{\text{nc}}$. Despite the application of NMS in conventional detectors, their direct $\text{mAP}_{\text{nc}}$ remains low due to numerous low-confidence false positives. To provide a fairer comparison, we perform threshold-based filtering to maximize $\text{mAP}_{\text{nc}}$ for conventional detectors. The figure clearly shows that incorporating confidence scores leads to a noticeable increase in mAP.
Instead of introducing an additional mechanism to estimate confidence for MLM-based detectors, we argue that confidence should not be a prerequisite when evaluating or comparing detection performance between MLMs and conventional detectors. Detection annotations and outputs inherently consist of class labels and bounding boxes, while confidence scores are auxiliary byproducts generated during inference. They may facilitate post-processing but are not indispensable for evaluating model accuracy. Therefore, we advocate employing confidence-independent metrics such as mean F1-score ($\text{mF}_{1}$) and $\text{mAP}_{\text{nc}}$ for a more equitable evaluation.
@@
Figure 5: The impact of confidence scores on $\text{mAP}_{\text{nc}}$ with error bands. The colored lines record the variation trends of $\text{mAP}_{\text{nc}}$ for the popular conventional detector on DOTA-v1.0 [dota] (trained and evaluated on both the ‘train‘ split and the ‘validation‘ split) dataset under different confidence thresholds.
To further verify the robustness of $\text{mAP}_{\text{nc}}$, we repeated each evaluation eleven times by substituting confidence values with ten random and one constant setting. As shown in Fig. 5, the resulting standard deviations remain below 0.5%, confirming its stability as a metric.
@@
Finally, for benchmarks such as DOTA [dota] and FAIR1M [FAIR1M], where public test sets are unavailable and online evaluation servers rely solely on mAP, we recommend adopting $\text{mAP}_{\text{nc}}$ as the primary evaluation metric to ensure consistent and fair assessment across different model types.
@@ IV Experiment
LLaVA-1.5 [llava15] 7B 0.1M 18.76 30.70 22.63 24.03
LLaVA-UHD-v2 [llavauhdv2] 7B 0.7M 22.82 32.57 26.08 27.16
Qwen2-VL [qwen2vl] 7B 11.1M 23.80 38.12 27.87 29.93
Qwen2.5-VL [qwen25vl] 7B 12.8M 19.66 35.82 26.70 27.39
Qwen3-VL [qwen25vl]
8B 16.8M 27.98 38.56 32.04 32.86
@@
InternVL2.5-MPO [mpo] 8B 2.4M 24.95 34.59 25.14 28.23
InternVL3 [internvl3] 8B 2.4M 22.49 38.09 26.36 28.98
InternVL3.5 [internvl35]
8B 2.4M 25.14 35.50 26.86 29.17
@@
Mimo-VL [xiaomi_mimovl] 7B 12.8M 16.51 20.04 27.11 21.22
GeoChat [geochat] 7B 0.3M 20.18 24.54 13.75 19.49
LLaVA-1.5 + SFT. on LRS-VQA [lrsvqa] 7B 0.1M 22.97 36.89 27.48 29.11
LLaVA-Next + SFT. on LRS-VQA [lrsvqa] 7B 2.8M 21.85 38.24 26.67 28.92
RSCoVLM
7B
1.0M
27.37 42.42 31.77 33.85
@@
Method
Input Size
DIOR-RSVG RSVG
GeoChat-VG
VRSBench-VG
AVVG
Avg. Acc.
@@
Qwen-VL-Chat [qwen2vl] $448\times 448$ 32.01 32.22 4.66 2.04 35.36 31.07 0.31 19.66
GeoChat [geochat] $504\times 504$ 23.35 24.05 3.08 2.04 22.74 11.52 0.28 12.44
LHRS-Bot [lhrsbot] $224\times 224$ 17.04 17.59 0.95 1.56 3.25 1.19 0.00 5.94
VHM [vhm] $336\times 336$ - 48.04 - - - - - -
Qwen2.5-VL [qwen25vl] Dynamic 43.64 45.26 19.73 21.27 42.99 44.50 7.64 32.15
Qwen-VL + SFT. on refGeo [geoground] $448\times 448$ 58.65 58.76 12.99 10.59 41.75 47.38 9.53 34.24
GeoChat + SFT. on refGeo [geoground] $504\times 504$ 60.27 61.96 16.32 14.67 56.99 51.36 11.52 39.01
LLaVA-1.5-7B + SFT. on refGeo [geoground] $336\times 336$ 64.46 65.98 19.98 20.95 63.76 57.17 15.05 43.91
GeoGround [geoground] $336\times 336$ 77.18 77.73 27.64 26.65 70.24 66.04 21.58 52.44
@@ IV-A Reproducibility Details
We have released the codebase on the  GitHub repository and uploaded the whole well-collected data folder and model weights to the [Uncaptioned image] HuggingFace repository. The codebase is implemented concisely, leveraging resource-efficient and effective training techniques. To save GPU memory, we adopt DeepSpeed-ZeRO-Stage-1 [zero] and gradient checkpointing. For improved computational efficiency, we utilize BFloat16 precision and Flash-Attention-2 [dao2024flashattention] during both training and evaluation. Additionally, Liger Kernel [hsu2025ligerkernel] is employed to accelerate training, and vLLM [vllm] is used for faster inferencing. All experiments are conducted on VolcEngine high-performance computing clusters equipped with NVIDIA A800 GPUs. We’ll maintain the repositories and update the latest code, model and data in our future research progress.
@@ IV-B1 Benchmark and Metric
There are three subsets, corresponding to three data sources: FAIR1M [FAIR1M], GLH-Bridge [li2024learning], and STAR [li2024scene]. The official scoring implementation first calculates accuracy for each source and task, and then computes average accuracy (AA) across tasks for each sources. The AAs for each subset are reported.
@@ IV-B2 Results
The results are presented in Table I. The max pixels numbers of each models are also provided. It can be seen that the average pixel number of LRS-VQA (about 45 million) has been larger than the largest pixels uplimit (16.8 million for Qwen3-VL [qwen25vl]).
@@
Furthermore, our model demonstrates stronger foundational capabilities than other competing models, approaching the performance of the leading Qwen3-VL-8B [qwen25vl], while utilizing a slightly smaller parameter count and a significantly lower maximum input resolution. Our model also outperforms other RS foundation models, including GeoChat and the officially fine-tuned LLaVA-Next model for LRS-VQA [lrsvqa].
@@ IV-C1 Benchmark and Metric
We follow GeoGround [geoground] for visual grounding evaluation because of its strong emphasis on comprehensiveness, fairness, and transparency. The evaluation incorporates the validation and test sets of DIOR-RSVG and RSVG [rsvg], the visual grounding portions of GeoChat-Bench [geochat] and VRSBench [vrsbench], as well as AVVG benchmark [geoground] for images captured by unmanned aerial vehicle. The evaluation details have strictly aligned with GeoGround. We directly adopted the splits and annotations provided by GeoGround for all benchmarks [geoground].
@@ IV-C2 Results
Table II presents the results, along with the corresponding input sizes for each model. The input resolutions of existing RS VLMs, including GeoChat [geochat], LHRS-Bot [lhrsbot], VHM [vhm], and GeoGround [geoground], are fixed and typically smaller than $\times512$. In contrast, only general-purpose VLMs such as Qwen2.5-VL [qwen25vl] and MiMo-VL [xiaomi_mimovl] support dynamic input resolution, enabling flexible adaptation to varying input sizes.
@@ IV-D Evaluation on Object Detection
GWD [gwd] 0.40 72.07 58.95 24.26 35.92 63.37 52.24 66.63 86.85 61.54 59.15 23.45 45.84 36.10 49.50 26.08 50.80 28.67 28.98
R3Det [r3det] 0.45 73.32 59.51 31.59 43.37 64.94 63.37 75.61 89.35 63.84 66.95 34.99 45.60 46.54 50.17 14.48 54.91 29.21 30.08
ATSS [atss] 0.35 72.98 60.67 25.82 42.91 65.23 65.32 75.22 89.78 71.61 70.12 28.04 43.19 47.79 58.28 28.60 56.37 34.62 33.05
Faster RCNN [fasterrcnn] 0.85 73.49 67.37 32.50 43.19 62.92 63.13 73.95 88.79 73.77 66.33 25.23 48.89 53.03 56.94 31.17 57.38 32.98 32.96
FCOS [fcos] 0.30 72.09 56.32 32.02 27.79 64.28 63.83 75.75 89.21 68.11 67.82 27.30 37.15 46.64 58.98 19.53 53.79 32.13 31.49
CSL [csl] 0.40 71.57 53.57 19.82 35.90 64.04 44.96 66.35 87.54 61.33 59.26 29.27 39.56 36.51 49.80 17.38 49.12 29.03 28.72
S2A-Net [s2anet] 0.50 72.96 61.96 36.00 45.99 66.24 65.61 77.08 89.34 73.27 69.25 31.78 46.64 55.02 52.02 35.40 58.58 29.53 31.98
@@
MiniGPTv2 [minigptv2] 7B 00.00|| 32.96- - 14.29 28.15 64.80
LLaVA-1.5 [llava15] 7B 00.00|| 31.10- - 21.73 34.96 54.55
Qwen-VL-Chat [qwen2vl] 7B 00.00|| 55.30- - 38.77 42.73 72.25
Qwen2.5-VL [qwen25vl] 7B 63.63 $|$ 62.73 70.90 56.64 64.98 76.20
Qwen3-VL [qwen25vl] 8B 70.84 $|$ 66.67 79.90 60.88 68.86 87.80
@@
InternVL2.5-MPO [mpo] 8B 69.38 $|$ 64.23 62.90 55.04 59.21 80.20
InternVL3 [internvl3] 8B 67.78 $|$ 63.40 67.29 59.65 64.32 86.40
InternVL3.5 [internvl35] 8B 77.03 $|$ 75.00 83.43 51.33 92.57 91.70
@@
Mimo-VL [xiaomi_mimovl] 7B 66.13 $|$ 67.20 69.14 54.51 64.35 86.10
LHRSBot [lhrsbot] 7B 00.00|| 91.26- - 69.81 83.94 93.17
GeoChat [geochat] 7B -72.00 $|$ -00.00 84.40 - - -
TEOChat [teochat] 7B -80.90 $|$ -00.00 86.30 - - -
LHRS-Bot-Nova [lhrsbot-nova] 7B -83.06 $|$ -00.00 - 72.74 83.97 96.20
SkysenseGPT [skysensegpt] 7B -88.16 $|$ -00.00 - 40.00 90.06 95.50
VHM [vhm] 7B 00.00|| 91.70- - 72.74 94.54 95.80
ScoreRS [scorers] 7B 00.00|| 85.90- - 74.42 91.59 96.30
@@
Method RSVQA Benchmark
VRSBench
VQA
@@
LLaVA-1.5 [llava15] 67.30 69.80 68.20 55.50 59.00 63.96 -
LLaVA-1.6 [scorers] 68.60 64.40 64.32 56.84 61.00 63.03 -
Qwen2-VL [qwen2vl] 75.60 63.30 75.47 62.00 73.00 69.87 -
Qwen2.5-VL [qwen25vl] 75.28 67.30 73.86 64.67 66.00 69.42 51.21
Qwen3-VL [qwen25vl] 81.00 78.10 70.32 56.42 72.00 71.57 54.75
InternVL-2.5 [internvl25] 75.50 65.80 71.16 66.21 72.00 70.13 47.20
InternVL3 [internvl3] 74.15 62.35 73.06 66.23 74.00 69.96 50.68
InternVL3.5 [internvl35] 80.14 53.80 92.00 91.26 96.00 82.64 53.74
Mimo-VL [xiaomi_mimovl] 66.00 77.80 74.42 59.98 65.00 68.64 48.37
LHRS-Bot-Nova [lhrsbot-nova] 89.30 87.60 88.11 83.89 79.00 85.58 -
GeoChat [geochat] 83.30 59.10 90.52 90.63 97.00 84.11 40.80
VHM [vhm] 83.30 68.30 90.11 89.89 87.00 83.72 -

### v2

@@ III-D Zoom-in Chain for UHR RS Images
Previous works on understanding UHR RS images, such as LRS-VQA [42] and GeoLLaVA-8k [55], primarily focus on addressing the issue of excessive image tokens through visual token pruning. Although this approach has proven effective and computationally efficient, it typically requires additional training and is not well-suited for joint training with tasks using standard image resolutions.
@@
$~~~~~~~~~~\texttt{Assistant}:~{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}[x1,y1,x2,y2]}$
@@
$~~~~~~~~~~\texttt{Assistant}:~{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{\textless Ground Truth\textgreater}}$ (5)
@@
To enable the model to learn zoom-in capabilities during training, we construct a specialized instruction tuning dataset for UHR RS image perception, named LRS-VQA-Zoom. The data pipeline is initiated by collecting three public, large-scale UHR RS datasets: DOTA1.0 [63], GLH-Bridge [31], and STAR [32].
The methodology for generating the LRS-VQA-Zoom is extended from the pipeline in LRS-VQA [42]. The final training corpus, totaling 302k samples, compromise three distinct subsets: 60k open-ended samples generated via rule-based templates, 159k open-ended samples synthesized using GPT-4V, and 83k samples in multi-choice-query form. Fig. 4 exhibits the examples from each subsets.
@@ III-E1 Response Normalization
During our extension of LMMRotate [27], we observed a subtle yet important issue. In the LMMRotate, images without any objects were removed from the training set to improve efficiency, following common practices in conventional aerial detector. However, this approach can be detrimental when training a VLM, as encountering object-free images during inference often leads the model to hallucinate, producing false positive detections. To address this, RSCoVLM retains images without objects in the training process and explicitly trains the model to output “There is none.” for such cases, thereby mitigating hallucinations and improving detection reliability.
@@ III-E2 Evaluation Metrics
In conventional aerial detection tasks, mean average precision (mAP) is widely employed as the evaluation metric, requiring bounding boxes, class labels, and confidence scores for all detected objects. However, as discussed earlier, our model responses only include object categories and their corresponding spatial coordinates, which implies that vision-language models based on the proposed detection approach cannot directly produce mAP results.
In mAP calculation, confidence scores are used to rank detector predictions, which directly influences the accumulation of true and false positives along the PR curve. The conventional object detection models often retain low-confidence yet actually false-positive predictions to maintain higher mAP scores. For instance, several DETR-based models evaluate all 900 proposals per image even when only a few objects are present. Since vision-language models typically yield only a limited number of confident predictions, we are constrained to assign fixed or randomized confidence scores to enable mAP computation. However, the lack of well-calibrated confidence estimates still places vision-language models at a substantial disadvantage in mAP-based evaluation with the adaptation, even when visual inspection suggests that VLMs achieve performance comparable to conventional detectors.
In practice, an object detection prediction is composed of a location and a category. The confidence scores are primarily used to filter out low-confidence detections via thresholding, as part of the detection result visualization procedure. If we first filter out low-confidence predictions using a threshold and then randomize the confidence scores, we could remove the influence of confidence on the mAP calculation. We denoted this as mAP with no conference scores, i.e., $\text{mAP}_{\text{nc}}$.
Figure 5 illustrates the $\text{mAP}_{\text{nc}}$ of several detectors. The horizontal axis illustrates the variation of $\text{mAP}_{\text{nc}}$ with increasing filtering thresholds. For each detector, the validation results were computed over ten runs with different random seeds, and the solid line represents the mean of these ten runs. The solid line in the figure is enveloped by a light-colored error band to indicate the effect of randomness on the results. Initially, as the filtering threshold increases, the detector’s performance improves because low-confidence false positives are progressively filtered out. After reaching a peak, overly strict filtering begins to remove true positives as well, causing the results to decline. Notably, this peak is still lower than the mAP metric that incorporates confidence scores. Furthermore, the error band is barely visible without magnification, indicating that although randomness is involved, the variance of $\text{mAP}_{\text{nc}}$ is very small.
@@
Figure 5: The impact of confidence scores on $\text{mAP}_{\text{nc}}$ with error bands. The colored lines record the variation trends of $\text{mAP}_{\text{nc}}$ for the popular conventional detector on DOTA-v1.0 [63] (trained and evaluated on both the ‘train‘ split and the ‘validation‘ split) dataset under different confidence thresholds.
Instead of introducing an additional mechanism to estimate confidence for VLM-based detectors, we argue that confidence should not be a prerequisite when evaluating or comparing detection performance between VLMs and conventional detectors. Detection annotations and outputs inherently consist of class labels and bounding boxes, while confidence scores are auxiliary byproducts generated during inference. They may facilitate postprocessing but are not indispensable for evaluating model accuracy. Therefore, we advocate employing confidence-independent metrics such as mean F1-score ($\text{mF}_{1}$) and $\text{mAP}_{\text{nc}}$ for a more equitable evaluation. Additionally, the small variance of $\text{mAP}_{\text{nc}}$ also demonstrates its stability as a metric.
@@
Finally, for benchmarks such as DOTA [63] and FAIR1M [51], where public test sets are unavailable and online evaluation servers rely solely on mAP, we recommend adopting $\text{mAP}_{\text{nc}}$ as the primary evaluation metric to ensure consistent and fair assessment across different model types.
@@ IV Experiment
LLaVA-1.5 [37] 7B 0.1M 18.76 30.70 22.63 24.03
LLaVA-UHD-v2 [75] 7B 0.7M 22.82 32.57 26.08 27.16
Qwen2-VL [56] 7B 11.1M 23.80 38.12 27.87 29.93
Qwen2.5-VL [2] 7B 12.8M 19.66 35.82 26.70 27.39
Qwen3-VL [2] 8B 16.8M 27.98 38.56 32.04 32.86
@@
InternVL2.5-MPO [58] 8B 2.4M 24.95 34.59 25.14 28.23
InternVL3 [82] 8B 2.4M 22.49 38.09 26.36 28.98
InternVL3.5 [59] 8B 2.4M 25.14 35.50 26.86 29.17
@@
Mimo-VL [70] 7B 12.8M 16.51 20.04 27.11 21.22
GeoChat [22] 7B 0.3M 20.18 24.54 13.75 19.49
LLaVA-1.5 + SFT. on LRS-VQA [42] 7B 0.1M 22.97 36.89 27.48 29.11
LLaVA-Next + SFT. on LRS-VQA [42] 7B 2.8M 21.85 38.24 26.67 28.92
RSCoVLM 7B 1.0M 27.37 42.42 31.77 33.85
@@
Method Input Size DIOR-RSVG RSVG GeoChat-VG VRSBench-VG AVVG Avg. Acc.
@@
Qwen-VL-Chat [56] $448\times 448$ 32.01 32.22 4.66 2.04 35.36 31.07 0.31 19.66
GeoChat [22] $504\times 504$ 23.35 24.05 3.08 2.04 22.74 11.52 0.28 12.44
LHRS-Bot [43] $224\times 224$ 17.04 17.59 0.95 1.56 3.25 1.19 0.00 5.94
VHM [46] $336\times 336$ - 48.04 - - - - - -
Qwen2.5-VL [2] Dynamic 43.64 45.26 19.73 21.27 42.99 44.50 7.64 32.15
Qwen-VL + SFT. on refGeo [78] $448\times 448$ 58.65 58.76 12.99 10.59 41.75 47.38 9.53 34.24
GeoChat + SFT. on refGeo [78] $504\times 504$ 60.27 61.96 16.32 14.67 56.99 51.36 11.52 39.01
LLaVA-1.5-7B + SFT. on refGeo [78] $336\times 336$ 64.46 65.98 19.98 20.95 63.76 57.17 15.05 43.91
GeoGround [78] $336\times 336$ 77.18 77.73 27.64 26.65 70.24 66.04 21.58 52.44
@@ IV-A Reproducibility Details
We have released the codebase on the  GitHub repository and uploaded the whole well-collected data folder and model weights to the [Uncaptioned image] HuggingFace repository. The codebase is implemented concisely, leveraging resource-efficient and effective training techniques. To save GPU memory, we adopt DeepSpeed-ZeRO-Stage-1 [48] and gradient checkpointing. For improved computational efficiency, we utilize BFloat16 precision and Flash-Attention-2 [14] during both training and evaluation. Additionally, Liger Kernel [19] is employed to accelerate training, and vLLM [23] is used for faster inferencing. All experiments are conducted on VolcEngine high-performance computing clusters equipped with NVIDIA A800 GPUs. We’ll maintain the repositories and update the latest code, model and data in our future research progress.
@@ IV-B1 Benchmark and Metric
There are three subsets, corresponding to three data sources: FAIR1M [51], GLH-Bridge [31], and STAR [32]. The official scoring implementation first calculates accuracy for each source and task, and then computes average accuracy (AA) across tasks for each sources. The AAs for each subset are reported.
@@ IV-B2 Results
The results are presented in Table I. The max pixels numbers of each models are also provided. It can be seen that the average pixel number of LRS-VQA (about 45 million) has been larger than the largest pixels uplimit (16.8 million for Qwen3-VL [2]).
@@
Furthermore, our model demonstrates stronger foundational capabilities than other competing models, approaching the performance of the leading Qwen3-VL-8B [2], while utilizing a slightly smaller parameter count and a significantly lower maximum input resolution. Our model also outperforms other RS foundation models, including GeoChat and the officially fine-tuned LLaVA-Next model for LRS-VQA [42].
@@ IV-C1 Benchmark and Metric
We follow GeoGround [78] for visual grounding evaluation because of its strong emphasis on comprehensiveness, fairness, and transparency. The evaluation incorporates the validation and test sets of DIOR-RSVG and RSVG [71], the visual grounding portions of GeoChat-Bench [22] and VRSBench [30], as well as AVVG benchmark [78] for images captured by unmanned aerial vehicle. The evaluation details have strictly aligned with GeoGround. We directly adopted the splits and annotations provided by GeoGround for all benchmarks [78].
@@ IV-C2 Results
Table II presents the results, along with the corresponding input sizes for each model. The input resolutions of existing RS VLMs, including GeoChat [22], LHRS-Bot [43], VHM [46], and GeoGround [78], are fixed and typically smaller than $\times512$. In contrast, only general-purpose VLMs such as Qwen2.5-VL [2] and MiMo-VL [70] support dynamic input resolution, enabling flexible adaptation to varying input sizes.
@@ IV-D Evaluation on Object Detection
GWD [67] 0.40 72.07 58.95 24.26 35.92 63.37 52.24 66.63 86.85 61.54 59.15 23.45 45.84 36.10 49.50 26.08 50.80 28.67 28.98
R3Det [66] 0.45 73.32 59.51 31.59 43.37 64.94 63.37 75.61 89.35 63.84 66.95 34.99 45.60 46.54 50.17 14.48 54.91 29.21 30.08
ATSS [73] 0.35 72.98 60.67 25.82 42.91 65.23 65.32 75.22 89.78 71.61 70.12 28.04 43.19 47.79 58.28 28.60 56.37 34.62 33.05
Faster RCNN [49] 0.85 73.49 67.37 32.50 43.19 62.92 63.13 73.95 88.79 73.77 66.33 25.23 48.89 53.03 56.94 31.17 57.38 32.98 32.96
FCOS [52] 0.30 72.09 56.32 32.02 27.79 64.28 63.83 75.75 89.21 68.11 67.82 27.30 37.15 46.64 58.98 19.53 53.79 32.13 31.49
CSL [68] 0.40 71.57 53.57 19.82 35.90 64.04 44.96 66.35 87.54 61.33 59.26 29.27 39.56 36.51 49.80 17.38 49.12 29.03 28.72
S2A-Net [17] 0.50 72.96 61.96 36.00 45.99 66.24 65.61 77.08 89.34 73.27 69.25 31.78 46.64 55.02 52.02 35.40 58.58 29.53 31.98
@@
MiniGPTv2 [6] 7B 00.00|| 32.96- - 14.29 28.15 64.80
LLaVA-1.5 [37] 7B 00.00|| 31.10- - 21.73 34.96 54.55
Qwen-VL-Chat [56] 7B 00.00|| 55.30- - 38.77 42.73 72.25
Qwen2.5-VL [2] 7B 63.63 $|$ 62.73 70.90 56.64 64.98 76.20
Qwen3-VL [2] 8B 70.84 $|$ 66.67 79.90 60.88 68.86 87.80
@@
InternVL2.5-MPO [58] 8B 69.38 $|$ 64.23 62.90 55.04 59.21 80.20
InternVL3 [82] 8B 67.78 $|$ 63.40 67.29 59.65 64.32 86.40
InternVL3.5 [59] 8B 77.03 $|$ 75.00 83.43 51.33 92.57 91.70
@@
Mimo-VL [70] 7B 66.13 $|$ 67.20 69.14 54.51 64.35 86.10
LHRSBot [43] 7B 00.00|| 91.26- - 69.81 83.94 93.17
GeoChat [22] 7B -72.00 $|$ -00.00 84.40 - - -
TEOChat [20] 7B -80.90 $|$ -00.00 86.30 - - -
LHRS-Bot-Nova [34] 7B -83.06 $|$ -00.00 - 72.74 83.97 96.20
SkysenseGPT [41] 7B -88.16 $|$ -00.00 - 40.00 90.06 95.50
VHM [46] 7B 00.00|| 91.70- - 72.74 94.54 95.80
ScoreRS [44] 7B 00.00|| 85.90- - 74.42 91.59 96.30
@@
Method RSVQA Benchmark VRSBench VQA
@@
LLaVA-1.5 [37] 67.30 69.80 68.20 55.50 59.00 63.96 -
LLaVA-1.6 [44] 68.60 64.40 64.32 56.84 61.00 63.03 -
Qwen2-VL [56] 75.60 63.30 75.47 62.00 73.00 69.87 -
Qwen2.5-VL [2] 75.28 67.30 73.86 64.67 66.00 69.42 51.21
Qwen3-VL [2] 81.00 78.10 70.32 56.42 72.00 71.57 54.75
InternVL-2.5 [8] 75.50 65.80 71.16 66.21 72.00 70.13 47.20
InternVL3 [82] 74.15 62.35 73.06 66.23 74.00 69.96 50.68
InternVL3.5 [59] 80.14 53.80 92.00 91.26 96.00 82.64 53.74
Mimo-VL [70] 66.00 77.80 74.42 59.98 65.00 68.64 48.37
LHRS-Bot-Nova [34] 89.30 87.60 88.11 83.89 79.00 85.58 -
GeoChat [22] 83.30 59.10 90.52 90.63 97.00 84.11 40.80
VHM [46] 83.30 68.30 90.11 89.89 87.00 83.72 -